The radius of the circle x^2+y^2-2x-2y+2=0
Thanks for the big help.. Kindly put the solution..

Answer by nerdybill(7384)

Standard equation of a circle is:
(x - h)^2 + (y - k)^2 = r^2
.
We "complete the square" to put our equation into the standard form:
x^2+y^2-2x-2y+2=0
x^2+y^2-2x-2y = -2
(x^2-2x)+(y^2-2y) = -2
(x^2-2x+1)+(y^2-2y+1) = -2+1+1
(x-1)^2+(y-1)^2 = 0
.
Radius = 0